How many grams of solid KCl are needed to prepare 250.0 mL of 0.235 M solution?
Answer #1

As, molarity = mass×1000/molar mass×volume(ml)

Molar mass of KCl = 74.55 g/mol

So, 0.235 = m×1000/74.55×250

m = 0.235×74.55×250/1000

= 4.38 g

Please give feedback!